# Neural Assets: 3D-Aware Multi-Object Scene Synthesis with Image Diffusion Models

### v1

@@ Abstract
We address the problem of multi-object 3D pose control in image diffusion models. Instead of conditioning on a sequence of text tokens, we propose to use a set of per-object representations, Neural Assets, to control the 3D pose of individual objects in a scene. Neural Assets are obtained by pooling visual representations of objects from a reference image, such as a frame in a video, and are trained to reconstruct the respective objects in a different image, e.g., a later frame in the video. Importantly, we encode object visuals from the reference image while conditioning on object poses from the target frame. This enables learning disentangled appearance and pose features. Combining visual and 3D pose representations in a sequence-of-tokens format allows us to keep the text-to-image architecture of existing models, with Neural Assets in place of text tokens. By fine-tuning a pre-trained text-to-image diffusion model with this information, our approach enables fine-grained 3D pose and placement control of individual objects in a scene. We further demonstrate that Neural Assets can be transferred and recomposed across different scenes. Our model achieves state-of-the-art multi-object editing results on both synthetic 3D scene datasets, as well as two real-world video datasets (Objectron, Waymo Open). Additional details and video results are available at our project page. Work done while interning at Google. Contact: tkipf@google.com
@@ 2 Related Work
2D spatial control in diffusion models (DMs). With the rapid growth of diffusion-based visual generation [90, 41, 67, 20, 75, 78], there have been many works aiming to inject spatial control to pre-trained DMs via 2D bounding boxes or segmentation masks. One line of research achieves this by manipulating text prompts [48, 30, 9], intermediate attention maps [104, 49, 15, 14, 25, 38, 10] or noisy latents [23, 86, 66, 61] in the diffusion process, without the need to change model weights. Closer to ours are methods that fine-tune pre-trained DMs to support additional spatial conditioning inputs [27, 3, 108, 42, 106, 31]. GLIGEN [55] introduces new attention layers to condition on bounding boxes. InstanceDiffusion [98] further supports object masks, points, and scribbles with a unified feature fusion block. To incorporate dense control signals such as depth maps and surface normals, ControlNet [110] adds zero-initialized convolution layers around the original network blocks. Recently, Boximator [95] demonstrates that such 2D control can be extended to video models with a similar technique. In our work, we build upon pre-trained DMs and leverage 3D bounding boxes as spatial conditioning, which enables 3D-aware control such as object rotation and occlusion handling.
3D-aware image generation. Earlier works leverage differentiable rendering to learn 3D Generative Adversarial Networks (GANs) [32] from monocular images, with explicit 3D representations such as radiance fields [12, 34, 13, 82, 68, 105] and meshes [16, 17, 29, 70, 71]. Inspired by the great success of DMs in image generation, several works try to lift the 2D knowledge to 3D [85, 59, 72, 94, 56, 46, 63, 100]. The pioneering work 3DiM [100] and follow-up work Zero-1-to-3 [57] directly train diffusion models on multi-view renderings of 3D assets. However, this line of research only considers single objects without background, which cannot handle in-the-wild data with complex backgrounds. Closest to ours are methods that process multi-object real-world scenes [80, 69, 109]. OBJect-3DIT [64] studies language-guided 3D-aware object editing by training on paired synthetic data, limiting its performance on real-world images [109]. LooseControl [6] converts 3D bounding boxes to depth maps to guide the object pose. Yet, it cannot be directly applied to edit existing images. In contrast, our Neural Asset representation captures both object appearance and 3D pose. It can be easily trained on real-world videos to achieve multi-object 3D edits.
@@ 3.3 Learning and Inference
Learning from frame pairs. As outlined in the introduction, we require a scalable data source of object-level "edits" in 3D space to effectively learn multi-object 3D control capabilities. Video data offers a natural solution to this problem: as the camera and the content of the scene moves or changes over time, objects are observed from various view points and thus in various poses and lighting conditions. We exploit this signal by randomly sampling pairs of frames from video clips, where we take one frame as the "source" image $x_{\mathrm{src}}$ and the other frame as the "target" image $x_{\mathrm{tgt}}$.
@@ 4.2 Main Results
Single-object editing. We first compare the ability to control the 3D pose of a single object on the OBJect dataset. Fig. 3 presents the results on the unseen object subset. We do not show FID here as it mainly measures the visual quality of generated examples, which does not reflect the editing accuracy. For results on the seen object subset and FID, please refer to Appendix B.1, where we observe similar trends. Compared to baselines, our model does not condition on text (e.g., the category name of the object to edit) as in 3DIT and is not pre-trained on multi-view rendering of 3D assets as in Zero-1-to-3. Still, we achieve state-of-the-art performance on all three tasks. This is because our Neural Assets representation learns disentangled appearance and pose features, which is able to preserve object identity while changing its placement smoothly. Also, the fine-tuned DINO encoder generalizes better to unseen objects compared to the frozen CLIP visual encoder used by baselines.
@@ 5 Conclusion
Limitations. An ideal Neural Asset should enable control over all potential configurations of an object such as deformation (e.g., a walking cat), rigid articulation (e.g., opening of a scissor), and structural decomposition (e.g., tomatoes being cut). In this work, we first tackle the foremost important aspect, i.e., controlling 3D rigid object pose and background composition which applies to almost all the objects. However, it can be adapted when suitable datasets are developed that capture other changes in objects. Another limitation is that our approach is currently limited to existing datasets that have 3D bounding box annotations. Yet, with recent advances in vision foundation models [52, 107, 7, 5, 58], we may soon have scalable 3D annotation pipelines similar to their 2D counterparts. However, this is out of scope for this work.
@@ Appendix D Broader Impacts
Potential negative societal impacts. Since we fine-tune large-scale pre-trained generative models in our pipeline, we inherit limitations of these base models, such as dataset selection bias. Such bias might be problematic when human subjects are involved, though our current approach is only capable of rigid object control and does not consider humans as an "asset" yet. Further study on how such bias affects model performance is required for mitigating negative societal impacts that could arise from this work.
Safeguards. For a potential future code release, we will be careful to apply automated content filters (e.g., NSFW classifiers) to the generation results of the model. There are no immediate plans to release model weights.

### v2

@@ Abstract
We address the problem of multi-object 3D pose control in image diffusion models. Instead of conditioning on a sequence of text tokens, we propose to use a set of per-object representations, Neural Assets, to control the 3D pose of individual objects in a scene. Neural Assets are obtained by pooling visual representations of objects from a reference image, such as a frame in a video, and are trained to reconstruct the respective objects in a different image, e.g., a later frame in the video. Importantly, we encode object visuals from the reference image while conditioning on object poses from the target frame. This enables learning disentangled appearance and pose features. Combining visual and 3D pose representations in a sequence-of-tokens format allows us to keep the text-to-image architecture of existing models, with Neural Assets in place of text tokens. By fine-tuning a pre-trained text-to-image diffusion model with this information, our approach enables fine-grained 3D pose and placement control of individual objects in a scene. We further demonstrate that Neural Assets can be transferred and recomposed across different scenes. Our model achieves state-of-the-art multi-object editing results on both synthetic 3D scene datasets, as well as two real-world video datasets (Objectron, Waymo Open). Additional details and video results are available at our project page. Work done while interning at Google.
Contact: tkipf@google.com. Project page: neural-assets.github.io
@@ 2 Related Work
2D spatial control in diffusion models (DMs). With the rapid growth of diffusion-based visual generation [94, 44, 70, 22, 79, 82], there have been many works aiming to inject spatial control to pre-trained DMs via 2D bounding boxes or segmentation masks. One line of research achieves this by manipulating text prompts [51, 32, 11], intermediate attention maps [110, 52, 17, 16, 27, 41, 12] or noisy latents [25, 90, 69, 64] in the diffusion process, without the need to change model weights. Closer to ours are methods that fine-tune pre-trained DMs to support additional spatial conditioning inputs [29, 5, 114, 45, 112, 33]. GLIGEN [59] introduces new attention layers to condition on bounding boxes. InstanceDiffusion [103] further supports object masks, points, and scribbles with a unified feature fusion block. To incorporate dense control signals such as depth maps and surface normals, ControlNet [116] adds zero-initialized convolution layers around the original network blocks. Recently, Boximator [99] demonstrates that such 2D control can be extended to video models with a similar technique. Several existing works [107, 3] leverage natural motion observed in video and similar to our work propose to train on paired video frames to achieve pixel-level control. In our work, we build upon pre-trained DMs and leverage 3D bounding boxes as spatial conditioning, which enables 3D-aware control such as object rotation and occlusion handling.
3D-aware image generation. Earlier works leverage differentiable rendering to learn 3D Generative Adversarial Networks (GANs) [34] from monocular images, with explicit 3D representations such as radiance fields [14, 37, 15, 86, 71, 111] and meshes [18, 19, 31, 73, 74]. Inspired by the great success of DMs in image generation, several works try to lift the 2D knowledge to 3D [89, 62, 75, 98, 60, 49, 66, 105]. The pioneering work 3DiM [105] and follow-up work Zero-1-to-3 [61] directly train diffusion models on multi-view renderings of 3D assets. However, this line of research only considers single objects without background, which cannot handle in-the-wild data with complex backgrounds. Closest to ours are methods that process multi-object real-world scenes [84, 72, 115, 2]. OBJect-3DIT [67] studies language-guided 3D-aware object editing by training on paired synthetic data, limiting its performance on real-world images [115]. LooseControl [8] converts 3D bounding boxes to depth maps to guide the object pose. Yet, it cannot be directly applied to edit existing images. In contrast, our Neural Asset representation captures both object appearance and 3D pose. It can be easily trained on real-world videos to achieve multi-object 3D edits.
From a methodology perspective, there have been prior works learning disentangled appearance and pose representations for 3D-aware multi-object image editing [100, 71, 111]. However, they are all based on the GAN framework [34] and do not learn generalizable object representations via an encoder. In contrast, we build upon a large-scale pre-trained image diffusion model [79] and powerful feature extractors [13], enabling editing of complex real-world scenes.
@@ 3.3 Learning and Inference
Learning from frame pairs. As outlined in the introduction, we require a scalable data source of object-level "edits" in 3D space to effectively learn multi-object 3D control capabilities. Video data offers a natural solution to this problem: as the camera and the content of the scene moves or changes over time, objects are observed from various view points and thus in various poses and lighting conditions [78]. We exploit this signal by randomly sampling pairs of frames from video clips, where we take one frame as the "source" image $x_{\mathrm{src}}$ and the other frame as the "target" image $x_{\mathrm{tgt}}$.
@@ 4.2 Main Results
Single-object editing. We first compare the ability to control the 3D pose of a single object on the OBJect dataset. Fig. 3 presents the results on the unseen object subset. We do not show FID here as it mainly measures the visual quality of generated examples, which does not reflect the editing accuracy. For results on the seen object subset and FID, please refer to Appendix B.1, where we observe similar trends. Compared to baselines, our model does not condition on text (e.g., the category name of the object to edit) as in 3DIT and is not trained on curated multi-view images of 3D assets as in Zero-1-to-3. Still, we achieve state-of-the-art performance on all three tasks. This is because our Neural Assets representation learns disentangled appearance and pose features, which is able to preserve object identity while changing its placement smoothly. Also, the fine-tuned DINO encoder generalizes better to unseen objects compared to the frozen CLIP visual encoder used by baselines.
@@ 4.4 Ablation Study
Figure 10: Failure case analysis. Our model mainly has two failure cases: (a) symmetry ambiguity, where the handle of the cup gets flipped when it rotates by 180 degrees; (b) camera-object motion entanglement, where the background also moves when we translate the foreground object. Both issues will likely be resolved if we train our Neural Assets model on more diverse data.
@@ 5 Conclusion
Limitations and future works. One main failure case of our model is symmetry ambiguity. As can be seen from the rotation results in Fig. 10 (a), the handle of the cup gets flipped when it rotates by 180 degree. Another failure case that only happens on Objectron is the entanglement of global camera motion and local object movement (Fig. 10 (b)). This is because Objectron videos only contain camera motion while objects always stay static. Both issues will likely be resolved if we train our model on larger-scale datasets with more diverse object and camera motion.
An ideal Neural Asset should enable control over all potential configurations of an object such as deformation (e.g., a walking cat), rigid articulation (e.g., opening of a scissor), and structural decomposition (e.g., tomatoes being cut). In this work, we first tackle the foremost important aspect, i.e., controlling 3D rigid object pose and background composition which applies to almost all the objects. Hence our current method does not allow for controlling structural changes. However, it can be adapted when suitable datasets are developed that capture other changes in objects.
Another limitation is that our approach is currently limited to existing datasets that have 3D bounding box annotations. Yet, with recent advances in vision foundation models [55, 113, 9], we may soon have scalable 3D annotation pipelines similar to their 2D counterparts. One notable example is OmniNOCS [56], which works on both Waymo and Objectron (datasets we used in this work), and diverse, in-the-wild Internet images for a wide range of object classes. It can be used to create larger open domain datasets to learn Neural Assets. We see this as an interesting future direction.
@@ Appendix D Broader Impacts
Potential negative societal impacts. Since we fine-tune large-scale pre-trained generative models in our pipeline, we inherit limitations of these base models, such as dataset selection bias. Such bias might be problematic when human subjects are involved, though our current approach is only capable of rigid object control and does not consider humans as an "asset" yet. Further study on how such bias affects model performance is required for mitigating negative societal impacts that could arise from this work [35].
Appendix E Funding Disclosure
This work was carried out at Google. Igor Gilitschenski contributed to the project in an advisory capacity.